OpenStudy (anonymous):

a jet ski company charges a flat fee of $20.00 plus $2.25 per hour to rent a jet ski. Another company charges a fee of $23.00 plus $2.75 per hour to rent the same jet ski. Using a graphing calculator, find the number of hours for which the costs are the same. Round your answer to the nearest whole hour. 1. 2 2. 3 3. 4 4. 8 I dont have a graphing calculator and my computer crashed when i use the online ones. Help!

OpenStudy (anonymous):

all i can tell ya is first company is 2.25x+20.00 and the second company is 2.75x+23.00

OpenStudy (anonymous):

so i have to graph it? I still don't know how to do it without a graphing calculator since its telling me to use one.

OpenStudy (anonymous):

i really dont understand how to solve the question :/

OpenStudy (texaschic101):

20 + 2.25h = 23 + 2..75h now solve for h(hours)

OpenStudy (texaschic101):

you have to set them equal to each other, then solve for h

OpenStudy (anonymous):

H = 3 so it is 3 hours correct?

OpenStudy (anonymous):

wait sorry I did something wrong!

OpenStudy (anonymous):

i would think

OpenStudy (texaschic101):

wait...there is something wrong...I am getting a negative number

OpenStudy (anonymous):

me too -6!

OpenStudy (texaschic101):

@mathstudent55 ...a little help here

OpenStudy (anonymous):

AH guys i put it wrong my fault. it's $24.00+ $2.25

OpenStudy (***[isuru]***):

Did u get the part @AllTheLonelyKillers said.. ? I mean first u have to find 2 functions 1st one... y = 2.25x+24.00 2 nd one y = 2.75x+23.00 and plot them in a single graph and the x co-ordinate intersection point of 2 lines is the answer.... since u don't have graphing calculator..... here's the link to the graph.. http://fooplot.com/plot/83jc9fzots hope this will help ya!!

OpenStudy (texaschic101):

thats better... 24 + 2.25h = 23 + 2.75h 24 - 23 = 2.75h - 2.25h 1 = .50h 1/.50 = h 2 = h

OpenStudy (anonymous):

yaaay i was partly right

OpenStudy (anonymous):

Thank you! Sorry about that haha now it makes sense

OpenStudy (texaschic101):

check... 24 + 2.25(2) = 23 + 2.75(2) 24 + 4.50 = 23 + 5.50 28.50 = 28.50 (correct)